OpenStudy (anonymous):

Simplify the following expression; express in terms of i. √ -49 √-12

OpenStudy (anonymous):

\[=-7\times 2\times \sqrt{3}\]\[=-14\sqrt{3}\]

OpenStudy (turingtest):

@patricia44 please use the equation feature to type your problem more clearly; copy-past is a bad idea usually

OpenStudy (anonymous):

simplify the following expression; express in termss of i. \[\sqrt -98\]

OpenStudy (turingtest):

first factor 98 completely